# UniRepLKNet: A Universal Perception Large-Kernel ConvNet for Audio, Video, Point Cloud, Time-Series and Image Recognition

### 27 Nov 2023 v1

@@ Abstract
Large-kernel convolutional neural networks (ConvNets) have recently received extensive research attention, but there are two unresolved and critical issues that demand further investigation. 1) The architectures of existing large-kernel ConvNets largely follow the design principles of conventional ConvNets or transformers, while the architectural design for large-kernel ConvNets remains under-addressed. 2) As transformers have dominated multiple modalities, it remains to be investigated whether ConvNets also have a strong universal perception ability in domains beyond vision. In this paper, we contribute from two aspects. 1) We propose four architectural guidelines for designing large-kernel ConvNets, the core of which is to exploit the essential characteristics of large kernels that distinguish them from small kernels - they can see wide without going deep. Following such guidelines, our proposed large-kernel ConvNet shows leading performance in image recognition. For example, our models achieve an ImageNet accuracy of 88.0%, ADE20K mIoU of 55.6%, and COCO box AP of 56.4%, demonstrating better performance and higher speed than a number of recently proposed powerful competitors. 2) We discover that large kernels are the key to unlocking the exceptional performance of ConvNets in domains where they were originally not proficient. With certain modality-related preprocessing approaches, the proposed model achieves state-of-the-art performance on time-series forecasting and audio recognition tasks even without modality-specific customization to the architecture. Code and all the models at https://github.com/AILab-CVC/UniRepLKNet.
@@ 1 Introduction
The design paradigm of convolutional neural networks (ConvNets) with very large kernels originated from RepLKNet [11] when the status of ConvNets was challenged by Vision Transformers (ViTs) [12, 38, 61, 65]. Inspired by ViTs that use global attention [12, 65, 54] or attention with large windows [38, 64, 49], RepLKNet proposed to use very large conv kernels. In contrast to the common practice using small kernels (e.g., 3$\times$3) [53, 19, 77, 22, 25, 58, 48], which fails to obtain a large Effective Receptive Field (ERF) [43] even with numerous small-kernel layers, RepLKNet realizes large ERF and impressive performance, especially on tasks such as object detection and semantic segmentation.
@@
We explore large-kernel ConvNet architecture by rethinking the design of conventional models that employ a deep stack of small kernels. As we add a 3$\times$3 conv to a small-kernel ConvNet, we expect it to take three effects simultaneously - 1) make the receptive field larger, 2) increase the abstract hierarchy of spatial patterns (e.g., from angles and textures to shapes of objects), and 3) improve the model’s general representational capability via making it deeper, bringing in more learnable parameters and non-linearities. In contrast, we argue that such three effects in a large-kernel architecture should be decoupled as the model should utilize the substantial strength of a large kernel - the ability to see wide without going deep. Since increasing the kernel size is much more effective than stacking more layers in enlarging the ERF [43], a sufficient ERF can be built up with a small number of large-kernel layers, so that the compute budget can be saved for other efficient structures that are more effective in increasing the abstract hierarchy of spatial patterns or generally increasing the depth. For example, when the objective is to extract higher-level local spatial patterns from lower-level ones, a 3$\times$3 conv layer might be a more suitable option than a large-kernel conv layer. The reason is that the latter demands more computations and may result in patterns that are no longer restricted to smaller local regions, which could be undesirable in specific scenarios.
@@
Our architecture achieves leading performance on ImageNet classification [9], ADE20K semantic segmentation [80], and COCO object detection [34], outperforming the existing large-kernel ConvNets such as RepLKNet [11], SLaK [36], and recent powerful architectures including ConvNeXt V2 [70], FastViT [63], Swin V2 [39] and DeiT III [62] in terms of both accuracy and efficiency.
RepLKNet [11] was proposed partly “in defense of ConvNets” as ViTs dominated multiple image recognition tasks that were once dominated by ConvNets. However, in this work, we seek to not only reclaim the leading position in image recognition tasks by surpassing ViTs’ performance but also demonstrate significant contributions in areas where ConvNets were not traditionally dominant. Specifically, on audio, video, point cloud, and time-series tasks, we achieve impressive performance with amazingly universal and simple solutions. We use modality-specific preprocessing approaches to transform all the data into 3D embedding maps just like what we do with images and use the same architecture as the backbone to process the embedding maps. Our model shows universal perception ability across multiple modalities with a unified architecture so it is named UniRepLKNet.
Impressively, UniRepLKNet achieves remarkable results even on modalities that were not considered the stronghold of ConvNet, such as audio and temporal data. For example, on a huge-scale time-series forecasting task that predicts the global temperature and wind speed, UniRepLKNet, which is a generalist model originally designed for image recognition, even outperforms the latest state-of-the-art transformer model customized for the task. Such results not only signify a comeback for ConvNet in its original domain but also showcase large-kernel ConvNet’s potential to conquer new territories, expanding its applicability and versatility in various tasks.
@@ 3.1 Dilated Reparam Block
Figure 2: Dilated Reparam Block uses dilated small-kernel conv layers to enhance a non-dilated large-kernel layer. Such dilated layers are equivalent to a non-dilated conv layer with a larger sparse kernel, as shown from the parameter perspective so that the whole block can be equivalently transformed into a single large-kernel conv. This example shows $K$=9 and we may use more dilated layers for larger $K$.
It is reported that a large-kernel conv should be used with a parallel small-kernel one because the latter helps capture the small-scale patterns during training [11]. Their outputs are added up after two respective Batch Normalization (BN) [28] layers. After training, with the Structural Re-parameterization [10] methodology, we merge the BN layers into the conv layers so that the small-kernel conv can be equivalently merged into the large-kernel one for inference. In this work, we note that except for small-scale patterns, enhancing the large kernel’s capability to capture sparse patterns (i.e., a pixel on a feature map may be more related to some distant pixels than its neighbors) may yield features of higher quality. The need to capture such patterns exactly matches the mechanism of dilated convolution - from a sliding-window perspective, a dilated conv layer with a dilation rate of $r$ scans the input channel to capture spatial patterns where each pixel of interest is $r-1$ pixels away from its neighbor. Therefore, we use dilated conv layers parallel to the large kernel and add up their outputs.
@@
Based on such equivalent transformations, we propose a Dilated Reparam Block, which uses a non-dilated small-kernel and multiple dilated small-kernel layers to enhance a non-dilated large-kernel conv layer. Its hyper-parameters include the size of large kernel $K$, sizes of parallel conv layers $\bm{k}$, and the dilation rates $\bm{r}$. The shown case (Fig. 2) with four parallel layers is denoted by $K$=9, $\bm{r}$=(1,2,3,4), $\bm{k}$=(5,3,3,3). For a larger $K$, we may use more dilated layers with larger kernel sizes or dilation rates. The kernel sizes and dilation rates of the parallel branches are flexible and the only constraint is $(k-1)r+1\leq K$. For example, with $K$=13 (the default setting in our experiments), we use five layers with $\bm{k}$=(5,7,3,3,3), $\bm{r}$=(1,2,3,4,5), so the equivalent kernel sizes will be (5,13,7,9,11), respectively. To convert a Dialted Reparam Block into a large-kernel conv layer for inference, we first merge every BN into the preceding conv layer, convert every layer with dilation $r>1$ with function 1, and add up all the resultant kernels with appropriate zero-paddings. For example, the layer in Fig. 2 with $k$=3,$r$=3 is converted into a sparse 7$\times$7 kernel and added to the 9$\times$9 kernel with one-pixel zero paddings on each side.
@@ 3.2 Architectural Guidelines for Large Kernels
Table 2: Models with different forms of Structural Re-parameterization on the 13$\times$13 conv layers.
@@
None N/A N/A 81.4 45.8
Dilated Reparam 5,7,3,3,3 1,2,3,4,5 81.6 46.5
Same kernel size 5,7,3,3,3 1,1,1,1,1 81.6 46.2
Same eq kernel size 5,13,7,9,11 1,1,1,1,1 81.6 46.2
@@
Architectural Guideline 4 on the Scaling Rule: while scaling up the depth, the added blocks should use small kernels. The scaling rule of existing large-kernel ConvNets follows the traditional ConvNets, i.e., stacking more large kernels to build up a deeper model, but we argue that a large-kernel ConvNet may not benefit from more large kernels. In this group of experiments (Table 4), we scale up $N_{3}$ from 9 to 27, following ConvNeXt-S [40]. Considering that nine 13$\times$13 blocks may have already built up sufficient receptive field, we examine if the added blocks should also use large kernels. Specifically, we refer to the block with a Dilated Reparam Block as the Large Kernel Block (LarK Block) and name a block that uses a DW 3$\times$3 conv as a Small Kernel Block (SmaK Block) so that the shallow model has 3 SmaK Blocks in Stage 1 and 3, 9, 3 LarK Blocks in the last three stages. While scaling up the depth of Stage 3, we tried the following options. A) All of the 27 blocks are LarK Blocks. B) We interleave SK with LarK Blocks so that Stage 3 has 14 LarK Blocks and 13 SmaK Blocks. C) We place two SmaK Blocks after a LarK Block so that the resultant model will have the same 9 LarK Blocks as before but 18 extra SmaK Blocks. D) We remove the DW 3$\times$3 layers in SmaK Blocks. Table 4 shows that scaling up the depth brings significant improvements, which is expected, and 9 LarK Blocks are sufficient. Though 27 LarK Blocks perform slightly better in the ADE20K mIoU, the inference speed is observably slowed down. Besides, the model without 3$\times$3 conv in SmaK Blocks shows significantly lower mIoU with only minor improvements in the throughput, suggesting such small kernels in SmaK Blocks are useful while scaling up the depth of large-kernel ConvNet as they increase the abstract hierarchy of spatial patterns, though they may not effectively enlarge the ERF [11, 43]. This observation supports our motivation to decouple the effects of conv layers in enlarging the ERF and extracting more complicated spatial patterns, as discussed in Sec. 1.
@@ 4 UniRepLKNet for Image Recognition
Table 8: Semantic segmentation on ADE20K validation set. The FLOPs are measured with 512$\times$2048 or 640$\times$2560 inputs according to the crop size. “SS” and “MS” mean single- and multi-scale testing, respectively. “${}^{\ddagger}$” indicates ImageNet-22K [9] pretraining.
@@ 5 Universal Perception on other Modalities
Time-series. Following Corrformer [71], we conduct experiments on the Global Temperature and Wind Speed Forecasting challenge  [^4] https://codeocean.com/capsule/0341365/tree/v1 using the dataset collected from the National Centers for Environmental Information (NCEI). This huge-scale dataset contains hourly averaged wind speed and temperature data from 3,850 stations with different geographical scales and densities, spanning from 2019 to 2021. For a fair comparison with Corrformer, which was the previous state-of-the-art method, we use its embedding layer (as introduced in Sec. 3.4) and decoder and only replace its encoder transformer with UniRepLKNet-S. We also compare UniRepLKNet-S against a wide range of methods including statistical and numerical approaches. We make the following observations from Table 9. 1) In terms of the forecasting precision, UniRepLKNet delivers a new state-of-the-art performance, achieving the lowest errors of 7.602, 1.832, 3.865, and 1.301 for MSE and MAE in forecasting global temperature and wind speed, respectively. 2) Regarding the efficiency, UniRepLKNet uses fewer parameters than existing deep learning methods. It is particularly noteworthy that UniRepLKNet, which is a generalist model, outperforms time-series specialists such as Pyraformer [35] and Corrformer [71] in both precision and efficiency. The significant advantages of UniRepLKNet open up new avenues for architectural discussions in time-series forecasting, presenting a viable alternative to transformer models.
Audio. We use Speech Commands V2 [69], which contains 105,829 one-second recordings of 35 common speech commands. Table 10 shows UniRepLKNet seamlessly adapts to the audio modality and delivers an impressive accuracy of 98.5%, even without pretraining. Compared to transformer-based methods such as AST [15] and Audio-MAE [26], UniRepLKNet stands out with fewer parameters. Compared to previous ConvNets designed for audio, UniRepLKNet achieves better performance without requiring any customizations to the structure, highlighting the untapped potential of ConvNets in the realm of audio recognition.
@@
Method Pretrain Type Acc (%) Params
@@
Video. We experiment on Kinetics-400 [29], which contains 240k training videos and 20k validation videos, spanning 400 classes for action recognition. Though its top-1 accuracy of 54.8% is somewhat behind state-of-the-art architectures like MViT [33] and TimeSFormer [2], we would like to note that the leading models are all specialized for video and pretrained with extra data, while UniRepLKNet is a generalist model trained from scratch. Compared to the latest generalist methods such as ImageBind [14] and Meta-Transformer [78], UniRepLKNet demonstrates superiority in both the accuracy and no requirement of pretraining. This observation highlights the challenge of achieving high transferability in the video domain with current general-purpose architectures, which may require a better preprocessing scheme than our extremely simple approach (laying out frames into a single image).
@@
Table 12: Point cloud analysis on ModelNet-40.
Method Type mAcc (%) OA (%)
@@
Impact of kernel size on the performance. To investigate the influence of different kernel sizes on performance, we replaced the backbone of the aforementioned models with other ConvNets or UniRepLKNet with smaller kernels. We adopted the same modality-specific preprocessing approaches and training configurations for a fair comparison. We take ResNet-101 as a representative small-kernel ConvNet because it has comparable parameters to UniRepLKNet-S. Table 13 shows that UniRepLKNet consistently outperforms other models by clear margins across all tasks, suggesting that the large-kernel design is crucial for universal perception, at least in our specific cases.

### 18 Mar 2024 v2

@@ Abstract
Large-kernel convolutional neural networks (ConvNets) have recently received extensive research attention, but two unresolved and critical issues demand further investigation. 1) The architectures of existing large-kernel ConvNets largely follow the design principles of conventional ConvNets or transformers, while the architectural design for large-kernel ConvNets remains under-addressed. 2) As transformers have dominated multiple modalities, it remains to be investigated whether ConvNets also have a strong universal perception ability in domains beyond vision. In this paper, we contribute from two aspects. 1) We propose four architectural guidelines for designing large-kernel ConvNets, the core of which is to exploit the essential characteristics of large kernels that distinguish them from small kernels - they can see wide without going deep. Following such guidelines, our proposed large-kernel ConvNet shows leading performance in image recognition (ImageNet accuracy of 88.0%, ADE20K mIoU of 55.6%, and COCO box AP of 56.4%), demonstrating better performance and higher speed than the recent powerful competitors. 2) We discover large kernels are the key to unlocking the exceptional performance of ConvNets in domains where they were originally not proficient. With certain modality-related preprocessing approaches, the proposed model achieves state-of-the-art performance on time-series forecasting and audio recognition tasks even without modality-specific customization to the architecture. All the code and models are publicly available on GitHub and Huggingface.
@@ 1 Introduction
The design paradigm of convolutional neural networks (ConvNets) with very large kernels originated from RepLKNet [19] when the status of ConvNets was challenged by Vision Transformers (ViTs) [20, 50, 74, 80, 22]. Inspired by ViTs that use global attention [20, 80, 67] or attention with large windows [50, 78, 62], RepLKNet proposed to use very large conv kernels. In contrast to the common practice using small kernels (e.g., 3$\times$3) [66, 28, 92, 31, 34, 71, 61], which fails to obtain a large Effective Receptive Field (ERF) [55] even with numerous small-kernel layers, RepLKNet realizes large ERF and impressive performance, especially on tasks such as object detection and semantic segmentation.
@@
We explore large-kernel ConvNet architecture by rethinking the design of conventional models that employ a deep stack of small kernels. As we add a 3$\times$3 conv to a small-kernel ConvNet, we expect it to take three effects simultaneously - 1) make the receptive field larger, 2) increase the abstract hierarchy of spatial patterns (e.g., from angles and textures to shapes of objects), and 3) improve the model’s general representational capability via making it deeper, bringing in more learnable parameters and non-linearities. In contrast, we argue that such three effects in a large-kernel architecture should be decoupled as the model should utilize the substantial strength of a large kernel - the ability to see wide without going deep. Since increasing the kernel size is much more effective than stacking more layers in enlarging the ERF [55], a sufficient ERF can be built up with a small number of large-kernel layers, so that the compute budget can be saved for other efficient structures that are more effective in increasing the abstract hierarchy of spatial patterns or generally increasing the depth. For example, when the objective is to extract higher-level local spatial patterns from lower-level ones, a 3$\times$3 might be a more suitable option than a large-kernel conv layer. The reason is that the latter demands more computations and may result in patterns no longer restricted to smaller local regions, which could be undesirable in specific scenarios.
@@
Our architecture achieves leading performance on ImageNet classification [12], ADE20K semantic segmentation [98], and COCO object detection [44], outperforming the existing large-kernel ConvNets such as RepLKNet [19], SLaK [48], and recent powerful architectures including ConvNeXt V2 [85], FastViT [77], Swin V2 [51] and DeiT III [75] in terms of both accuracy and efficiency. Moreover, our architecture demonstrates significantly higher shape bias [76, 3] than existing ConvNets and ViTs, i.e., it makes predictions more based on the overall shapes of objects than the textures, which agrees with the human visual system and results in better generalization. This may explain its superiority in downstream tasks. See the Appendix for details.
RepLKNet [19] was proposed partly “in defense of ConvNets” as ViTs dominated multiple image recognition tasks that were once dominated by ConvNets. Moreover, considering transformers have shown universal perception capability in multiple modalities [94, 93], in this work, we seek to not only reclaim the leading position in image recognition tasks by surpassing ViTs’ performance but also contribute to areas where ConvNets were not traditionally dominant. Specifically, on audio, video, point cloud, and time-series tasks, we achieve impressive performance with amazingly universal and simple solutions. We use modality-specific preprocessing approaches to transform all the data into 3D embedding maps just like what we do with images and use the same architecture as the backbone to process the embedding maps. Our model shows universal perception ability across multiple modalities with a unified architecture so it is named UniRepLKNet.
Impressively, UniRepLKNet achieves remarkable results even on modalities that were not considered the stronghold of ConvNet, e.g., audio and temporal data. On a huge-scale time-series forecasting task that predicts the global temperature and wind speed, UniRepLKNet, a generalist model originally designed for image recognition, even outperforms the latest state-of-the-art transformer customized for the task. Such results not only signify a comeback for ConvNet in its original domain but also showcase large-kernel ConvNet’s potential to conquer new territories, expanding its applicability and versatility in various tasks.
@@ 3 Architectural Design of UniRepLKNet
We first summarize the architectural guidelines as follows. 1) Block design: use efficient structures that perform both inter-channel communications and spatial aggregations to increase the depth. 2) Re-parameterization: use dilated small kernels to re-parameterize a large kernel. 3) Kernel size: decide kernel size according to the downstream task and usually use large kernels in middle- and high-level layers. 4) Scaling Rule: while scaling up the depth, the added blocks should use small kernels. We describe the proposed Dilated Reparam Block in Sec. 3.1 and details in Sec. 3.2.
@@ 3.1 Dilated Reparam Block
Figure 2: Dilated Reparam Block uses dilated small-kernel conv layers to enhance a non-dilated large-kernel layer. Such dilated layers are equivalent to a non-dilated conv layer with a larger sparse kernel, as shown from the parameter perspective, so that the whole block can be equivalently transformed into a single large-kernel conv. This example shows $K$=9, and we may use more dilated layers for larger $K$.
It is reported a large-kernel conv should be used with a parallel small-kernel one because the latter helps capture the small-scale patterns during training [19]. Their outputs are added up after two respective Batch Normalization (BN) [37] layers. After training, with the Structural Re-parameterization [17, 13, 16, 14, 15, 18] methodology, we merge the BN layers into the conv layers so the small-kernel conv can be equivalently merged into the large-kernel one for inference. In this work, we note that except for small-scale patterns, enhancing the large kernel’s capability to capture sparse patterns (i.e., a pixel on a feature map may be more related to some distant pixels than its neighbors) may yield features of higher quality. The need to capture such patterns exactly matches the mechanism of dilated convolution - from a sliding-window perspective, a dilated conv layer with a dilation rate of $r$ scans the input channel to capture spatial patterns where each pixel of interest is $r-1$ pixels away from its neighbor. Therefore, we use dilated conv layers parallel to the large kernel and add up their outputs.
@@
Based on such equivalent transformations, we propose a Dilated Reparam Block, which uses a non-dilated small-kernel and multiple dilated small-kernel layers to enhance a non-dilated large-kernel conv layer. Its hyper-parameters include the size of large kernel $K$, sizes of parallel conv layers $\bm{k}$, and the dilation rates $\bm{r}$. The shown case (Fig. 2) with four parallel layers is denoted by $K$=9, $\bm{r}$=(1,2,3,4), $\bm{k}$=(5,5,3,3). For a larger $K$, we may use more dilated layers with larger kernel sizes or dilation rates. The kernel sizes and dilation rates of the parallel branches are flexible and the only constraint is $(k-1)r+1\leq K$. For example, with $K$=13 (the default setting in our experiments), we use five layers with $\bm{k}$=(5,7,3,3,3), $\bm{r}$=(1,2,3,4,5), so the equivalent kernel sizes will be (5,13,7,9,11), respectively. To convert a Dialted Reparam Block into a large-kernel conv layer for inference, we first merge every BN into the preceding conv layer, convert every layer with dilation $r>1$ with function 1, and add up all the resultant kernels with appropriate zero-paddings. For example, the layer in Fig. 2 with $k$=3,$r$=3 is converted into a sparse 7$\times$7 kernel and added to the 9$\times$9 kernel with one-pixel zero paddings on each side.
@@ 3.2 Architectural Guidelines for Large Kernels
Table 2: Different forms of Structural Re-parameterization on the 13$\times$13 conv layers. We report the mean±$\pm$±std of three runs.
@@
None N/A N/A 81.44$\pm$0.04 45.78$\pm$0.05
Dilated Reparam 5,7,3,3,3 1,2,3,4,5 81.63$\pm$0.02 46.37$\pm$0.10
Same kernel size 5,7,3,3,3 1,1,1,1,1 81.55$\pm$0.01 46.07$\pm$0.07
Same eq kernel size 5,13,7,9,11 1,1,1,1,1 81.59$\pm$0.02 46.17$\pm$0.04
@@
Architectural Guideline 4 on the Scaling Rule: while scaling up the depth, the added blocks should use small kernels. The scaling rule of existing large-kernel ConvNets follows the traditional ConvNets, i.e., stacking more large kernels to build up a deeper model, but we argue that a large-kernel ConvNet may not benefit from more large kernels. In this group of experiments (Table 4), we scale up $N_{3}$ from 9 to 27, following ConvNeXt-S [52]. Considering that nine 13$\times$13 blocks may have already built up sufficient receptive field, we examine if the added blocks should also use large kernels. Specifically, we refer to the block with a Dilated Reparam Block as the Large Kernel Block (LarK Block) and name a block that uses a DW 3$\times$3 conv as a Small Kernel Block (SmaK Block) so that there are 3 SmaK Blocks in Stage 1 and 3/9/3 LarK Blocks in Stage 2/3/4 of the shallow model. While scaling up the depth of Stage 3, we tried the following options. A) All of the 27 blocks are LarK Blocks. B) We interleave SmaK with LarK Blocks so that Stage 3 has 14 LarK Blocks and 13 SmaK Blocks. C) We place two SmaK Blocks after a LarK Block so that the resultant model will have the same 9 LarK Blocks as before but 18 extra SmaK Blocks. D) We remove the DW 3$\times$3 layers in SmaK Blocks. Table 4 shows that scaling up the depth brings significant improvements, which is expected, and 9 LarK Blocks are sufficient. Though 27 LarK Blocks perform slightly better in the ADE20K mIoU, the inference speed is observably slowed down. Besides, the model without 3$\times$3 conv in SmaK Blocks shows significantly lower mIoU with only minor improvements in the throughput, suggesting such small kernels in SmaK Blocks are useful while scaling up the depth of large-kernel ConvNet as they increase the abstract hierarchy of spatial patterns, though they may not effectively enlarge the ERF [19, 55]. This observation supports our motivation to decouple the effects of conv layers in enlarging the ERF and extracting more complicated spatial patterns, as discussed in Sec. 1.
@@ 4 UniRepLKNet for Image Recognition
Table 8: Semantic segmentation on ADE20K validation set. The FLOPs are measured with 512$\times$2048 or 640$\times$2560 inputs according to the crop size. “SS” and “MS” mean single- and multi-scale testing, respectively. “${}^{\ddagger}$” ImageNet-22K [12] pretraining.
@@ 5 Universal Perception on other Modalities
Time-series. Following Corrformer [86], we conduct experiments on the Global Temperature and Wind Speed Forecasting challenge  [^4] https://codeocean.com/capsule/0341365/tree/v1 using the dataset collected from the National Centers for Environmental Information (NCEI). This huge-scale dataset contains hourly averaged wind speed and temperature data from 3,850 stations with different geographical scales and densities, spanning from 2019 to 2021. For a fair comparison with Corrformer, which was the previous state-of-the-art method, we use its embedding layer (as introduced in Sec. 3.4) and decoder and only replace its encoder transformer with UniRepLKNet-S. We also compare UniRepLKNet-S against a wide range of methods, including statistical and numerical approaches. Table 9 shows UniRepLKNet delivers a new state-of-the-art forecasting precision, achieving the lowest errors of 7.602, 1.832, 3.865, and 1.301 for MSE and MAE in forecasting global temperature and wind speed, respectively, with fewer parameters than existing deep learning methods. It is particularly noteworthy that UniRepLKNet, a generalist model, outperforms time-series specialists such as Pyraformer [47] and Corrformer [86] in both precision and efficiency. The significant advantages of UniRepLKNet open up new avenues for architectural discussions in time-series forecasting, presenting a viable alternative to transformer models.
Audio. We use Speech Commands V2 [84], which contains 105,829 one-second recordings of 35 common speech commands. Table 10 shows UniRepLKNet seamlessly adapts to audio and delivers an impressive accuracy of 98.5%, even without pretraining. Compared to transformers such as AST [24] and Audio-MAE [35], UniRepLKNet stands out with fewer parameters. Compared to previous ConvNets designed for audio, UniRepLKNet achieves better performance without customizations to the structure, highlighting the untapped potential of ConvNets in the realm of audio.
@@
Method Pretrain Type Acc. (%) Params
@@
Video. Kinetics-400 [39] contains 240k training videos and 20k validation videos, spanning 400 classes for action recognition. Though the top-1 accuracy of 54.8% is somewhat behind state-of-the-art architectures like MViT [43], we note that UniRepLKNet is a generalist model without pretraining. Compared to the latest generalist methods, ImageBind [23] and Meta-Transformer [93], UniRepLKNet shows higher accuracy and requires no pretraining.
@@
Table 12: Point cloud analysis on ModelNet-40 dataset.
Method Type ModelNet-40
mAcc (%) OA (%)
@@
Impact of kernel size on the performance. To investigate the influence of different kernel sizes on performance, we compare UniRepLKNet with models of smaller kernels. We adopted the same modality-specific preprocessing approaches and training configurations for a fair comparison. We take ResNet-101 as a representative small-kernel ConvNet because it has comparable parameters to UniRepLKNet-S. Table 13 shows large kernels are crucial for universal perception, at least in our specific cases.